OpenStudy (anonymous):

Protons are accelerated by electromagnets around a circular chamber with a radius of 1 km to speeds near the speed of light before colliding with a target to produce large amounts of energy. If a proton is traveling at 10% the speed of light, how much centripetal force is exerted by the electromagnets? (speed of light is 3 X 10^8 m/s and the mass of a proton is 1.67 X 10^-27 kg)

OpenStudy (abb0t):

Use: \[F_c = \frac{ mv^2 }{ r }\] m= mass of one proton v= 10% of speed of light r= radius (IN METERS)

OpenStudy (anonymous):

I know to use that formula but I am having trouble putting the numbers into it. The mass is 1.67*10^-27

OpenStudy (abb0t):

You already have the given information. Just plug it in.

OpenStudy (abb0t):

You might want to convert 1 km -> metres.

OpenStudy (anonymous):

Which would be .001m?

OpenStudy (abb0t):

no.

OpenStudy (abb0t):

1000 metres

OpenStudy (anonymous):

I figured it out its 1000m

OpenStudy (anonymous):

Our teacher gave us the answer which is 1.5 *10^-15N, but no matter what I do i just cant get this number

OpenStudy (abb0t):

Maybe your teacher is wrong then. But it should be ~1.5E-13

OpenStudy (anonymous):

What did you do to get the answer

OpenStudy (abb0t):

same thing ur teacher got, except to the -13 instead of -15.

OpenStudy (anonymous):

In the formula I plug in 1.67*10^-27 for the mass, 10 for the speed and 1000 for the radius but the answer i get is 1.67*10^-29. I don't know what I'm doing wrong

OpenStudy (abb0t):

It must be what you're typing in your calculator because I get 1.5...

OpenStudy (anonymous):

Do I have the correct numbers above?

OpenStudy (abb0t):

no

OpenStudy (anonymous):

Is the radius correct? 1000m

OpenStudy (abb0t):

yes.

OpenStudy (anonymous):

Would V equal 6283.18

OpenStudy (abb0t):

\[\frac{ (1.67 \times 10^{-27})(3 \times 10^8)^2 }{ 1000 }\]

OpenStudy (anonymous):

I probably was forgetting the parenthesis

OpenStudy (anonymous):

I put the parenthesis and got the answer.